OpenStudy (anonymous):

reciprocal of 1and 1/2

OpenStudy (amistre64):

\[\frac{1}{1~and~1/2}\]

OpenStudy (amistre64):

\[\frac{1}{1+1/2}\] \[\frac{2(1)}{2(1+1/2)}\] \[\frac{2}{2+1}\]

OpenStudy (cggurumanjunath):